OpenStudy (anonymous):

MEDALS!!! If f(x) = 0.6(3 – x), what is the value of f(–3)? A. 0 B. 1.8 C. 3.6 D. 4.8

jimthompson5910 (jim_thompson5910):

Everywhere you see an x, replace it with -3. Then evaluate

jimthompson5910 (jim_thompson5910):

f(x) = 0.6(3 – x) f(-3) = 0.6(3 – (-3)) f(-3) = ???

geerky42 (geerky42):

\(f(\color{red}x) = 0.6 (3-\color{red}x) \\ f(\color{red}{-3}) = 0.6 (3-(\color{red}{-3}))\)

geerky42 (geerky42):

you just replace x to whether what x was replaced to in f(x)

geerky42 (geerky42):

So you simply replace x to -3, then solve it.

OpenStudy (anonymous):

So I got C 3.6

geerky42 (geerky42):

right!